OpenStudy (anonymous):

solve 65/n=5/8 please show work answer should be {104}

OpenStudy (anonymous):

65/n = 5/8 n = 65*8/5 = 13*8 = 104

OpenStudy (anonymous):

what does * mean

OpenStudy (roadjester):

65/n = 5/8 cross multiply 8*65=5n then divide by 5 to solve for n n=104

OpenStudy (anonymous):

multiplication.

OpenStudy (roadjester):

* means multiply

OpenStudy (roadjester):

x is usually used as a variable so that's why * is used instead

OpenStudy (anonymous):

thanks you guys